OpenStudy (anonymous):

A right circular cylinder has a base whose area is 616 square feet. If the height of the cylinder is 9 feet, what is its lateral surface area? Use π = 22/7

OpenStudy (anonymous):

all you need to do is find the radius of the base of the cylinder. solve for r: 616 = pi*r^2 then stick the value you got for r into this to get the lateral surface area: S = 2pi*r*h = 2(22/7)*9*r